# A Correspondence-Driven Approach for Bilevel Decision-making with Nonconvex Lower-Level Problems

### v1

@@ 4.2 Geometric Analysis near Bifurcation Points
Intuitively, the Minkowski dimension reflects how the number of small balls needed to cover the set grows as the ball radius shrinks. For a smooth embedded submanifold, it matches the topological dimension.
@@
Suppose $g(x,y)$ is a semi-algebraic function for $(x,y)$, and Assumption 2.1 holds, then the Minkowski dimension $d$ of $\widetilde{\mathcal{X}}$ is less than or equal to n1n-1.
@@
The core of the proof of Theorem 4.1 is that the bifurcation set $\widetilde{\mathcal{X}}$ corresponding to a semi-algebraic function $g(x,y)$ is itself a semi-algebraic set, and thus admits a stratified manifold structure. Since $\widetilde{\mathcal{X}}$ has Lebesgue measure zero, all manifolds in the stratification must have dimension at most n1n-1, which implies that the Minkowski dimension of $\widetilde{\mathcal{X}}$ is at most n1n-1. A concrete example of such a stratified manifold structure for $\widetilde{\mathcal{X}}$ is illustrated in Figure 5. The same structure extends to the real-analytic case, but for simplicity, we do not introduce tools from subanalytic geometry.
Figure 5: Visualization of the bifurcation point set $\mathcal{X}$ of the function $g(x,y)=y^{4}+(x_{1}^{2}-5x_{1}x_{2}+2x_{2}^{2}-7x_{1}+8x_{2}-30)y^{3}+(x_{1}^{2}-3x_{1}x_{2}+4x_{2}^{2}-5x_{1}+2x_{2}-40)y^{2}+(x_{1}^{2}-5x_{1}x_{2}+2x_{2}^{2}-7x_{1}+8x_{2}-30)y.$ over the domain $4,5]^{2}$. The set exhibits a finite stratified manifold structure: red points correspond to 0-dimensional manifolds, while blue curves represent 1-dimensional manifolds.
@@
The estimate on the measure of $\widetilde{\mathcal{X}}_{\delta}$ allows us to control the first term in (4.7), which is essential for establishing Proposition 4.1. As a key intermediate result, this proposition is indispensable for the proof of Theorem 4.2.
$\delta$$\delta$$2\delta$$\tilde{\mathcal{X}}$
@@ 4.3 Upper-level Oracle Complexity Analysis
We now turn to the second term in the integral decomposition (4.7), which corresponds to the region $\mathbb{R}^{n}/\widetilde{\mathcal{X}}_{\delta}.$ In this region, we aim to control the approximation error $\|y^{\text{cd}}(x)-\hat{y}(x)\|,$ where $y^{\text{cd}}(x)$ is the exact solution defined in (3.1), and $\hat{y}(x)$ is the output of the cubic-regularized Newton method (see Algorithm 2) after finite $K$ steps which independent of $x$.
@@
At a high level, our approach to analyzing the algorithm convergence is as follows: since we focus on points $x$ outside the bifurcation set $\widetilde{\mathcal{X}}$, $g(x,y)$ is Morse in $y$, and all of its stationary points are isolated and non-degenerate. At each such point, the norm of every Hessian eigenvalue is bounded below by $\mu>0$ (to be shown in Lemma 4.2). Thus, by the Lipschitz continuity of Hessian from Assumption 3.2(4), within a open ball of radius $0<r\leq\mu/(2\overline{\overline{L}}_{g})$ centered at any stationary point in $\mathcal{Y}$, all eigenvalues of $\nabla^{2}_{yy}g(x,y)$ have absolute values lower bounded by $\mu/2$. As illustrated in Figure 7, we denote these open neighborhoods by blue-shaded disks centered at each stationary point. Outside the union of all radius-$r$ open balls centered at stationary points, the remaining points in $\mathcal{Y}$ form a compact set that contains no stationary point. By continuity of $\nabla_{y}g(x,y)$, it follows that $\|\nabla_{y}g(x,y)\|$ is bounded below by a positive constant on this compact set. These structural conditions allow us to characterize the iterates produced by the cubic-regularized Newton method and to derive non-asymptotic bounds on the approximation error $\|\hat{y}(x)-y^{\mathrm{cd}}(x)\|$. Specifically, known results for cubic Newton (Nesterov and Polyak, 2006) guarantee that after $K$ steps the algorithm produces a point $\hat{y}$ such that
@@
To formalize this high level idea, we now state two technical lemmas. Lemma 4.2 guarantees that outside a small $\delta$-neighborhood of the bifurcation set $\widetilde{\mathcal{X}}$, every stationary point $y\in\mathcal{Y}^{\prime}$ of $g(x,\cdot)$ has a Hessian $\nabla^{2}_{y}g(x,y)$ whose eigenvalues are uniformly bounded away from zero. Here $\mathcal{Y}^{\prime}$ is a slightly enlarged compact superset of $\mathcal{Y}$, introduced for technical convenience in the proof. Lemma 4.3 guarantees that for any $x\in\mathcal{X}\setminus\widetilde{\mathcal{X}}_{\delta}$, the gradient norm $\|\nabla_{y}g(x,y)\|$ is uniformly lower bounded by a positive constant $\alpha(\delta,r)$ for all $y$ away from the $r$-neighborhood of the stationary set. The constant $\alpha(\delta,r)$ depends only on $r$ and $\delta$, not on $x$. These two results play a key role in characterizing the behavior of the cubic-regularized Newton method.
@@
Suppose Assumption 3.2 holds. Let $\mathcal{Y}^{\prime}\supset\mathcal{Y}$ be a compact set with $\mathcal{Y}\subset\mathrm{int}(\mathcal{Y}^{\prime})$, where $\mathcal{Y}$ is defined in Assumption 3.2(6). Then there exists a positive function $\mu(\delta)$ which only depends on $\delta$ such that for any $x\in\mathcal{X}\setminus\widetilde{\mathcal{X}}_{\delta}$ and any $y\in\mathcal{Y}^{\prime}$ that is a stationary point of $g(x,\cdot)$, all eigenvalues $\lambda$ of $\nabla^{2}_{yy}g(x,y)$ satisfy $|\lambda|\geq\mu(\delta)$.
@@
where $\overline{g}_{0}$ and $\underline{g}$ are from Proposition 3.1(1) and Assumption 3.2(3), functions $\mu(\cdot)$ and $\alpha(\cdot,\cdot)$ are from Lemma 4.2 and Lemma 4.3, and $\rho>0$ is a constant which represent the approximation error in the lower-level solution, then we have
@@
Suppose Assumptions 3.1-3.44.1 hold, and iteration number of inner loop $K$ satisfies (4.4) depends on $\delta$. Define $\nabla^{K}_{x}\phi^{\text{cd}}_{\xi}(x)$ as follows:
@@
for some constant $C_{1}(n,\xi)$, $C_{2}(n,\xi)$ which depend on $n$ and $\xi$, where $d$ is the Minkowski dimension of $\widetilde{\mathcal{X}}$ assumed in Assumpetion 4.1.
@@
Combining the above bias and variance bounds, we are now in a position to analyze the convergence behavior of the proposed algorithm. The statement and proof can be viewed as a variant of Lan (2020, Theorem6.6): in Lan (2020, Theorem6.6), the analysis is carried out for mirror descent methods, whereas here we specialize Lan (2020, Equation (6.2.1)) to the case $h\equiv 0$ and take $V$ in Lan (2020, Equation (6.2.6)) to be the squared Euclidean distance, which yields projected SGD. The key difference from Lan (2020, Theorem 6.6) is that our gradient estimator is biased, while theirs is unbiased.
@@
$K_{t}>\max\left\{\frac{256(\overline{g}_{0}-\underline{g})\left(\overline{\overline{L}}_{g}\right)^{1/2}}{9\left(\min\left\{\alpha(\delta_{t},\frac{\mu_{g}(\delta_{t})}{2\overline{\overline{L}}_{g}}),\frac{(\mu_{g}(\delta_{t}))^{2}}{4\overline{\overline{L}}_{g}}\right\}\right)^{3/2}},\frac{768(\overline{g}_{0}-\underline{g})\overline{\overline{L}}^{2}_{g}}{\left(\mu_{g}(\delta_{t})\right)^{3}}\right\}+\frac{256(\overline{g}_{0}-\underline{g})\left(\overline{\overline{L}}_{g}\right)^{1/2}}{9(\mu_{g}(\delta_{t})\rho_{t})^{3/2}}.$ (4.15)
@@ 4.4 Lower-level Oracle Complexity under Fold Bifurcation Assumption
The second condition in Definition 4.3 requires that a tiny change of the parameter in some direction instantly changed the stationarity condition along the degenerate direction $v$. It is a prevalent condition, with prevalence understood in the parametric-family sense discussed in Remark 2.1. Specifically, it is not hard to see that the second condition is equivalent to requiring that the matrix
@@
It is easy to compute that the stationary points of $g(x,y)$ with respect to $y$ near $y=0$ is
@@
Suppose $g(x,y)$ is smooth, and all degenerate stationary points of $g(x,y)$ with respect to $y$ are fold bifurcation points defined in Definition 4.3, then there exist constants $D_{1},D_{2},D_{3},D_{4}>0$ such that for any $\delta>0$
@@
Under the setting of Theorem 4.2, we assume that $g(x,y)$ is smooth, and all degenerate stationary points of $g(x,y)$ with respect to $y$ are fold bifurcation points (Definition 4.3). Let $\rho_{t}=1/(t+1)$, $\delta_{t}=(t+1)^{-2/(n-d)}$, $N_{t}=t+1$ and use a constant step size $\beta_{t}=\beta<\overline{L}_{\phi^{\text{cd}}_{\xi}}/2$. To achieve $\mathbb{E}\big{[}\|\widetilde{\Phi}_{\mathcal{X},R}\|^{2}\big{]}\leq\epsilon^{2}$, we have the following oracle complexity:
@@ 5.1 Numerical Example
The function $f(x,y)$ is neither convex in $x$ nor concave in $y$. We simulate the GDA flow using 15 initializations with random seeds from 0 to 14, and observe that 4 trajectories (with seeds 5, 10, 12, and 13) enter closed loops, indicating non-convergence. We apply our SCiNBiO to this nonconvex-nonconcave problem, and set the algorithmic parameters as follows:

### v3

@@ 4.2 Geometric Analysis near Bifurcation Points
Intuitively, the Minkowski dimension reflects how the number of small balls needed to cover the set grows as the ball radius shrinks.
@@
Suppose $g(x,y)$ is a semi-algebraic function for $(x,y)$, and Assumption 2.1 holds, then the upper Minkowski dimension $d$ of $\widetilde{\mathcal{X}}$ is less than or equal to n1n-1.
@@
The core of the proof of Theorem 4.1 is that the bifurcation set $\widetilde{\mathcal{X}}$ corresponding to a semi-algebraic function $g(x,y)$ is itself a semi-algebraic set, and thus admits a stratified manifold structure. Since $\widetilde{\mathcal{X}}$ has Lebesgue measure zero, all manifolds in the stratification must have dimension at most n1n-1, which implies that the Minkowski dimension of $\widetilde{\mathcal{X}}$ is at most n1n-1. A concrete example of such a stratified manifold structure for $\widetilde{\mathcal{X}}$ is illustrated in Figure 5.
Figure 5: Visualization of the bifurcation point set $\widetilde{\mathcal{X}}$ of the function $g(x,y)=y^{4}+(x_{1}^{2}-5x_{1}x_{2}+2x_{2}^{2}-7x_{1}+8x_{2}-30)y^{3}+(x_{1}^{2}-3x_{1}x_{2}+4x_{2}^{2}-5x_{1}+2x_{2}-40)y^{2}+(x_{1}^{2}-5x_{1}x_{2}+2x_{2}^{2}-7x_{1}+8x_{2}-30)y.$ over the domain $4,5]^{2}$. The set exhibits a finite stratified manifold structure.
@@
The estimate on the measure of $\widetilde{\mathcal{X}}_{\delta}$ allows us to control the first term in (4.7), which is essential for establishing Proposition 4.1. As a key intermediate result, this proposition is indispensable for the proof of Theorem 4.2, which establishes the complexity of Algorithm (1) for the upper-level problem.
$\delta$$\delta$$2\delta$$\widetilde{\mathcal{X}}$
@@ 4.3 Upper-level Oracle Complexity Analysis
We now turn to the second term in the integral decomposition (4.7), which corresponds to the region $\mathbb{R}^{n}\setminus\widetilde{\mathcal{X}}_{\delta}.$ In this region, we aim to control the approximation error $\|y^{\text{cd}}(x)-\hat{y}(x)\|,$ where $y^{\text{cd}}(x)$ is the exact solution defined in (3.1), and $\hat{y}(x)$ is the output of the cubic-regularized Newton method (see Algorithm 2) after finite $K$ steps which are independent of $x$.
@@
At a high level, our approach to analyzing the algorithm convergence is as follows: since we focus on points $x$ outside the bifurcation set $\widetilde{\mathcal{X}}$, $g(x,y)$ is Morse in $y$, and all of its stationary points are isolated and non-degenerate. At each such point, the norm of every Hessian eigenvalue is bounded below by $\mu>0$ (to be shown in Lemma 4.2). Thus, by the Lipschitz continuity of Hessian from Assumption 3.2(4), within an open ball of radius $0<r\leq\mu/(2\overline{\overline{L}}_{g})$ centered at any stationary point in $\mathcal{Y}$, all eigenvalues of $\nabla^{2}_{yy}g(x,y)$ have absolute values lower bounded by $\mu/2$. As illustrated in Figure 7, we denote these open neighborhoods by blue-shaded disks centered at each stationary point. Outside the union of all radius-$r$ open balls centered at stationary points, the remaining points in $\mathcal{Y}$ form a compact set that contains no stationary point. By continuity of $\nabla_{y}g(x,y)$, it follows that $\|\nabla_{y}g(x,y)\|$ is bounded below by a positive constant on this compact set. These structural conditions allow us to characterize the iterates produced by the cubic-regularized Newton method and to derive non-asymptotic bounds on the approximation error $\|\hat{y}(x)-y^{\mathrm{cd}}(x)\|$. Specifically, known results for cubic Newton (Nesterov and Polyak, 2006) guarantee that after $K$ steps the algorithm produces a point $\hat{y}$ such that
@@
To formalize this high level idea, we now state two technical lemmas. Lemma 4.2 guarantees that outside $\widetilde{\mathcal{X}}_{\delta}$, which is $\delta$-neighborhood of the bifurcation set $\widetilde{\mathcal{X}}$, every stationary point $y\in\mathcal{Y}^{\prime}$ of $g(x,\cdot)$ has a Hessian $\nabla^{2}_{y}g(x,y)$ whose eigenvalues are uniformly bounded away from zero. Here $\mathcal{Y}^{\prime}$ is a slightly enlarged compact superset of $\mathcal{Y}$, introduced for technical convenience in the proof. Lemma 4.3 guarantees that for any $x\in\mathcal{X}\setminus\widetilde{\mathcal{X}}_{\delta}$, the gradient norm $\|\nabla_{y}g(x,y)\|$ is uniformly lower bounded by a positive constant $\alpha(\delta,r)$ for all $y$ away from the $r$-neighborhood of the stationary set. The constant $\alpha(\delta,r)$ depends only on $r$ and $\delta$, not on $x$. These two results play a key role in characterizing the behavior of the cubic-regularized Newton method.
@@
Suppose Assumption 3.2 holds. Let $\mathcal{Y}^{\prime}\supset\mathcal{Y}$ be a compact set with $\mathcal{Y}\subset\mathrm{int}(\mathcal{Y}^{\prime})$, where $\mathcal{Y}$ is defined in Assumption 3.2(6). Then there exists a positive function $\mu(\delta)$ depending only on $\delta$ such that for any $x\in\mathcal{X}\setminus\widetilde{\mathcal{X}}_{\delta}$ and any stationary points $y\in\mathcal{Y}^{\prime}$ of $g(x,\cdot)$, all eigenvalues $\lambda$ of $\nabla^{2}_{yy}g(x,y)$ satisfy $|\lambda|\geq\mu(\delta)$.
@@
where $\overline{g}_{0}$ and $\underline{g}$ are from Proposition 3.1(1) and Assumption 3.2(3), functions $\mu(\cdot)$ and $\alpha(\cdot,\cdot)$ are from Lemma 4.2 and Lemma 4.3, and $\rho>0$ is a constant which represents the approximation error in the lower-level solution, then we have
@@
Suppose Assumptions 3.1-3.44.1 hold, and the iteration number of the inner loop $K$ satisfies (4.4) depends on $\delta$. Define $\nabla^{K}_{x}\phi^{\text{cd}}_{\xi}(x)$ as follows:
@@
for some constant $C_{1}(n,\xi)$, $C_{2}(n,\xi)$ which depend on $n$ and $\xi$, where $d$ is the Minkowski dimension of $\widetilde{\mathcal{X}}$ assumed in Assumption 4.1.
@@
With the bias and variance bounds in place, we proceed to analyze the convergence of the proposed algorithm. The statement and proof of the following Theorem 4.2 can be viewed as a variant of Lan (2020, Theorem6.6): in Lan (2020, Theorem6.6), the analysis is carried out for mirror descent methods, whereas here we specialize Lan (2020, Equation (6.2.1)) to the case $h\equiv 0$ and take $V$ in Lan (2020, Equation (6.2.6)) to be the squared Euclidean distance, which yields projected SGD. The key difference from Lan (2020, Theorem 6.6) is that our gradient estimator is biased, while theirs is unbiased.
@@
$K_{t}>\max\left\{\frac{256(\overline{g}_{0}-\underline{g})\left(\overline{\overline{L}}_{g}\right)^{1/2}}{9\left(\min\left\{\alpha(\delta_{t},\frac{\mu(\delta_{t})}{2\overline{\overline{L}}_{g}}),\frac{(\mu(\delta_{t}))^{2}}{4\overline{\overline{L}}_{g}}\right\}\right)^{3/2}},\frac{768(\overline{g}_{0}-\underline{g})\overline{\overline{L}}^{2}_{g}}{\left(\mu(\delta_{t})\right)^{3}}\right\}+\frac{256(\overline{g}_{0}-\underline{g})\left(\overline{\overline{L}}_{g}\right)^{1/2}}{9(\mu(\delta_{t})\rho_{t})^{3/2}}.$ (4.15)
@@ 4.4 Lower-level Oracle Complexity under Fold Bifurcation Assumption
The second condition in Definition 4.3 requires that a tiny change of the parameter in some direction instantly changes the stationarity condition along the degenerate direction $v$. It is a prevalent condition, with prevalence understood in the parametric-family sense discussed in Remark 2.1. Specifically, it is not hard to see that the second condition is equivalent to requiring that the matrix
@@
It is easy to compute that the stationary points of $g(x,y)$ with respect to $y$ near $y=0$ are
@@
Suppose $g(x,y)$ is three times continuously differentiable, and all degenerate stationary points of $g(x,y)$ with respect to $y$ are fold bifurcation points defined in Definition 4.3, then there exist constants $D_{1},D_{2},D_{3},D_{4}>0$ such that for any $\delta>0$
@@
Under the setting of Theorem 4.2, we assume that $g(x,y)$ is three times continuously differentiable, and all degenerate stationary points of $g(x,y)$ with respect to $y$ are fold bifurcation points (Definition 4.3). Let $\rho_{t}=1/(t+1)$, $\delta_{t}=(t+1)^{-2/(n-d)}$, $N_{t}=t+1$ and use a constant step size $\beta_{t}=\beta<\overline{L}_{\phi^{\text{cd}}_{\xi}}/2$. To achieve $\mathbb{E}\big[\|\widetilde{\Phi}_{\mathcal{X},R}\|^{2}\big]\leq\epsilon^{2}$, we have the following oracle complexity:
@@ 5.1 Numerical Example
The function $f(x,y)$ is neither convex in $x$ nor concave in $y$. We simulate the GDA flow using 15 initializations with random seeds from 0 to 14, and observe that 4 trajectories (with seeds 5, 10, 12, and 13) enter closed loops, indicating non-convergence. We apply our algorithm SCiNBiO to this nonconvex-nonconcave problem, and set the algorithmic parameters as follows:
@@
5.2 Hyperparameter Optimization
In this section, we evaluate the performance of SCiNBiO against V-PBGD (Shen et al., 2025) and BOME (Liu et al., 2022a) on a bilevel hyperparameter optimization task. The goal is to tune a nonnegative $\ell_{2}$-regularization parameter $\lambda$ for a neural network.
We model the problem as choosing $\lambda$ to minimize the validation loss, subject to the network weights $\theta$ being optimized on the training set. To enforce $\lambda\geq 0$, we parameterize it as $\lambda=\mathrm{softplus}(x)=\log(1+e^{x})$ with $x\in\mathbb{R}$. Formally, we solve:
$\min_{\lambda\geq 0}\quad f(\theta^{\ast}(\lambda),\lambda)$ $:=\mathrm{CE}_{\mathrm{val}}(\theta^{\ast}(\lambda))$ (5.2a)
$\text{s.t.}\quad\theta^{\ast}(\lambda)\in\mathop{\mathrm{argmin}}_{\theta}\quad g(\theta,\lambda)$ $:=\mathrm{CE}_{\mathrm{train}}(\theta)+\tfrac{\lambda}{2B}\|\theta\|_{2}^{2},$ (5.2b)
where $\mathrm{CE}{\mathrm{train}}$ and $\mathrm{CE}{\mathrm{val}}$ denote the cross-entropy loss on training and validation sets, respectively, and $B$ is the training batch size. The cross-entropy loss for a dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$ is defined as:
$\mathrm{CE}(\theta;{(x_{i},y_{i})}_{i=1}^{n})=-\frac{1}{n}\sum_{i=1}^{n}\log\left(\frac{\exp(z_{\theta}(x_{i}){y_{i}})}{\sum_{j=1}^{10}\exp(z_{\theta}(x_{i})j)}\right),$ (5.3)
where $z_{\theta}(x_{i})$ represents the classifier output.
We use the MNIST dataset with a split of $B=50000$ training examples for the lower-level problem (5.2b) and $10000$ validation examples for the upper-level objective (5.2a). The classifier $z_{\theta}(\cdot)$ is a Convolutional Neural Network (CNN). To ensure the mapping $\theta\mapsto z_{\theta}(\cdot)$ remains differentiable, we employ SiLU activations and average pooling. The specific architecture is:
[…]
This deep architecture creates a highly nonconvex lower-level landscape.
We apply SCiNBiO (our proposed Algorithm 1), V-PBGD (Shen et al., 2025), and BOME (Liu et al., 2022a) to the bilevel hyperparameter optimization (5.2a). The specific experimental settings and algorithmic parameters are as follows: for all three algorithms, the lower-level optimization is performed using SGD with a step size $\eta=0.1$, a batch size of $B=2048$, and $K=100$ iterations per inner solve. The additional parameters for each algorithm are specified below:
  • SCiNBiO: outer loop epoch number $T=100$, warm-started inner solves (i.e., initialize each inner-loop run at the final iterate from the previous inner-loop run, rather than restarting from the same fixed $y_{0}$ in Algorithm 1), outer loop step size $\beta_{t}=0.001/\sqrt{t}$ using square-root decay, number of samples $N_{\mathrm{dirs}}=3$ for upper-level gradient estimator, smoothing radius $\xi=0.1$;
  • V-PBGD: outer loop epoch number $T=300$, outer step size $\beta=0.01$, penalty parameter $\rho_{t}=\min\{1.05^{t-1},10\}$;
  • BOME: outer loop epoch number $T=300$, outer loop step size $\eta=0.1$.
To evaluate the performance, we measure both the upper-level validation loss and the lower-level objective, tracking the convergence over epochs. In Figure 12, we present the performance of SCiNBiO, V-PBGD, and BOME on this hyperparameter optimization task, evaluated over 30 random seeds. The left plot and the right plot represent the upper-level validation loss and the lower-level objective, respectively, both plotted against the number of epochs. Since SCiNBiO performs $N_{\mathrm{dirs}}=3$ inner-loop cycles for each outer-loop iteration, the plots are shown against the number of inner-loop epochs to provide a consistent comparison across the algorithms. The deep-colored curves represent the average performance across the 30 seeds for each algorithm, while the light-colored curves show the individual results from each of the 30 seeds. The results, shown in Figure 12, highlight the superior performance of SCiNBiO. As seen in the figure, SCiNBiO consistently achieves faster convergence and more stable optimization compared to V-PBGD and BOME, with strong performance in both the upper-level and lower-level tasks.
Figure 12: Plot of the upper-level & lower-level function vs epoch for seeds 0 to 29